OpenStudy (anonymous):

area of a circle of diamter d is 100 square inches. The area of another circle of radius d is 200 square inches 100 square inches 400 square inches 24 square inhes 100 and 24 are obviously out of the question whats the correct answer? Also how did you find the answer?

OpenStudy (faiqraees):

pi (d/2)^2=100 d^2=400/pi Now substitue the value of d^2 in the formula of area of second circle and you will get the answer

OpenStudy (anonymous):

thank you!